# MIT-CTP/5902

### v1

@@ 4.2 Identifying $\Pi$ that flows to $\Theta$
To show that the symmetry generated by $\Pi$ and $\mathcal{Q}$ is anomaly-free, consider the following Hamiltonian, which is symmetric and trivially gapped: [^24] Following footnote 22, one could ask how we identified $\Pi$ as the operator that flows to $\Theta$, rather than $\Pi_{\text{e}}=\mathcal{X}_{\text{even}}\mathcal{R}\mathcal{T}=X\Pi$, which flows to $\exp(i\pi\mathsf{Q}_{\text{w}})\Theta$. The key point is that the lattice anomaly between $\mathcal{X}$ and $\mathcal{Q}$ leads to a mixed anomaly between $\Pi_{\text{e}}$ and $\mathcal{Q}$. See the discussion in Subsection 4.5.1. On the other hand, there is no anomaly between $\Pi$ and $\mathcal{Q}$ as we show here.
@@ 4.3 Self-anomaly of $\mathcal{U}$
which also probes the self-anomaly of $\mathcal{U}$.
@@ 4.4 Anomaly for $\tilde{\Pi}$}
Given that there is a mixed anomaly between $\Pi$ and $\mathcal{U}$, it is natural to ask if $\tilde{\Pi}$} itself is anomalous. We will follow the discussion in Section 3.3.1 to identify a mod 2 anomaly for $\tilde{\Pi}$}. We will restrict ourselves to the even $L$ case, and separate the discussion into two cases: $L=0,2$ mod 4.252525In the discussion in Section 4.1, we limited ourselves to $L=0\bmod 4$. This was necessary for the properties of $\mathcal{Q}$. However, here we do not discuss $\mathcal{Q}$, so we can study the model for any even $L$. It will become clear that these two cases are analogous to the even and odd $L$ cases in Section 3.3.1.
@@
${\tilde{\Pi}}^{2}=-1\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ L=2\leavevmode\nobreak\ \text{mod}\leavevmode\nobreak 4\,$, (4.32)
@@ 4.5.1 The lattice anomaly between $\mathcal{X}$ and $e^{i\pi\mathcal{Q}}$
Here, we use $\Pi$ to probe the emergent anomaly between $e^{i\frac{\pi}{2}\mathsf{Q}_{\mathrm{w}}}$ and $e^{i\pi\mathsf{Q}_{\mathrm{m}}}$.
@@ 5 Non-invertible RT
    $\mathcal{D}X_{j}=Z_{j}Z_{j+1}\mathcal{D}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{D}Z_{j}Z_{j+1}=X_{j+1}\mathcal{D}\,.$ (5.3)
@@
    $\mathcal{T}_{\text{Ising}}X_{j}\mathcal{T}_{\text{Ising}}^{-1}=X_{j}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{T}_{\text{Ising}}Y_{j}\mathcal{T}_{\text{Ising}}^{-1}=-Y_{j}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{T}_{\text{Ising}}Z_{j}\mathcal{T}_{\text{Ising}}^{-1}=Z_{j}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{T}_{\text{Ising}}i\mathcal{T}_{\text{Ising}}^{-1}=-i\,.$ (5.6)
@@
${\tilde{\Pi}}_{\text{Ising}}X_{j}=Z_{-j}Z_{-j+1}{\tilde{\Pi}}_{\text{Ising}}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\tilde{\Pi}}_{\text{Ising}}Z_{j}Z_{j+1}=X_{-j}{\tilde{\Pi}}_{\text{Ising}}\,.$ (5.8)
@@
$({\tilde{\Pi}}_{\text{Ising}})^{2}=1+\eta\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\tilde{\Pi}}_{\text{Ising}}\eta=\eta{\tilde{\Pi}}_{\text{Ising}}={\tilde{\Pi}}_{\text{Ising}}\,,\quad\eta^{2}=1\,.$ (5.9)
@@
$i\chi_{2j-1}\chi_{2j}=X_{j}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ i\chi_{2j}\chi_{2j+1}=Z_{j}Z_{j+1}\,.$ (5.11)
@@ Acknowledgments
We are grateful to Maissam Barkeshli, Boris Bulatovic, Meng Cheng, Lukasz Fidkowski, Tarun Grover, Ryohei Kobayashi, Ho Tat Lam, Max Metlitski, Greg Moore, Salvatore Pace, Abhinav Prem, Sahand Seifnashri, Nikita Sopenko, Xincheng Zhang, and Yifan Zhang for interesting discussions. The work of NS was supported in part by DOE grant DE-SC0009988. SHS was supported in part by NSF grant PHY-2449936. WZ was supported in part by the Princeton University Department of Physics. This work was also supported by the Simons Collaboration on Ultra-Quantum Matter, which is a grant from the Simons Foundation (651444, NS, SHS).
@@ A.1 Non-spin theories
Let us apply this reasoning to a field theory. As mentioned in footnote 29, now we should use $\mathscr{R}\to\mathsf{R}$, $\mathscr{T}\to\mathsf{T}$, and $\varTheta\to\Theta$. In addition to the invariance under $\Theta$, the theory might also have additional symmetries. These can be internal symmetries, which do not act on the spacetime coordinate, as well as spacetime symmetries like space-reflection and time-reversal. We would like to make a few comments about the consequences of that.
@@ A.2 Spin theories
In Euclidean signature, we lift $SO(d+1)$ to $Spin(d+1)$ and again $R_{ij}(2\pi)=(-1)^{F}$ with $i,j=1,\cdots,d+1$ and $i\neq j$. [^30] Even though we do not focus on the individual reflection elements $\mathscr{R}^{i}$, but only on their bilinears $\mathscr{R}^{i}\mathscr{R}^{j}=R_{ij}(\pi)$, we note that $R_{ij}(\pi)=(-1)^{F}R_{ji}(\pi)$. And therefore, $\mathscr{R}^{i}\mathscr{R}^{j}=(-1)^{F}\mathscr{R}^{j}\mathscr{R}^{i}$. (This is regardless of whether $(\mathscr{R}^{i})^{2}$ is $1$ or $(-1)^{F}$, i.e., regardless of whether they extend $Spin(d+1)$ to $Pin^{+}(d+1)$ or to $Pin^{-}(d+1)$.) This means that after the lift $R_{ij}(\pi)^{2}=(-1)^{F}$, including
@@
Rotating this back to Lorentzian signature, $\mathscr{R}^{d+1}$ becomes the anti-unitary transformation $\mathscr{T}$. Therefore, $\mathscr{R}\mathscr{R}^{d+1}$ becomes the anti-unitary transformation $\varTheta=\mathscr{R}\mathscr{T}$ and its square has another factor of $(-1)^{F}$ to give [^31] As in footnote 30, $\mathscr{R}\mathscr{T}=(-1)^{F}\mathscr{T}\mathscr{R}$, and using (A.16), $\mathscr{R}^{2}=(-1)^{F}\mathscr{T}^{2}$. (Again, this is regardless of whether $\mathscr{R}^{2}$ is $1$ or $(-1)^{F}$.)
@@
This crucial fact will be discussed in detail below for the special case of $d=1$. The extension to $d>1$ is straightforward.
@@
We claim that a Hermitian operator $\cal O$ with spin $s$ has [^33] Importantly, in order for the eigenvalues of anti-unitary operators to be well-defined, we let them act on Hermitian operators. This fact is at the root of the $\varTheta$ assignment for spinors in (LABEL:Rof_O).
@@
In order to understand the significance of the assignment of $\varTheta({\cal O})$, let us discuss the multiplication of these representations. Consider two Hermitian operators, ${\cal O}_{i}$ with spins $s_{i}$ and $\varTheta$ and $(-1)^{F}$ as in (LABEL:Rof_O). If $s_{1}$ is an integer, ${\cal O}_{1}{\cal O}_{2}$ is Hermitian. Its total spin is $s_{1}+s_{2}$ and $\varTheta({\cal O}_{1}{\cal O}_{2})=(-1)^{s_{1}}(-1)^{s_{2}(2s_{2}-1)}=(-1)^{(s_{1}+s_{2})(2s_{1}+2s_{2}-1)}$. The same conclusion applies if $s_{2}$ is an integer. However, if both $s_{1}$ and $s_{2}$ are half-integers, these two operators are fermions. Then, to determine the $\varTheta$-eigenvalue of their product, we should consider the Hermitian operator $i{\cal O}_{1}{\cal O}_{2}$. (See footnote 33.) Its spin is $s_{1}+s_{2}$ and $\varTheta(i{\cal O}_{1}{\cal O}_{2})=-(-1)^{s_{1}(2s_{1}-1)}(-1)^{s_{2}(2s_{2}-1)}=(-1)^{s_{1}+s_{2}}$. (This reasoning provides a derivation of the $\varTheta$-eigenvalue assignments in (LABEL:Rof_O) by multiplying representations with $s=\pm{1\over 2}$.)
@@
Repeating the discussion about odd spatial dimensions in (A.13), we have
@@ Appendix B Review of time-reversal anomalies in quantum mechanics of fermions
$[(-1)^{F}]^{2}=1\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{T}^{2}=(-1)^{\frac{\nu(\nu-2)}{8}}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{T}(-1)^{F}=(-1)^{\frac{\nu}{2}}(-1)^{F}\mathcal{T}\,,\quad\nu=0\bmod 2\,.$ (B.10)
@@
$\mathcal{T}^{\prime 2}=(-1)^{\frac{\nu(\nu+2)}{8}}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{T}^{\prime}(-1)^{F}=(-1)^{\frac{\nu}{2}}(-1)^{F}\mathcal{T}^{\prime}\,,\quad\nu=0\bmod 2\,.$ (B.11)
@@
We consider $N_{f}$ non-chiral, massless, free Majorana fermion fields in the continuum (2.16), with their left- and right-moving components denoted by $\chi_{\text{L,R}}^{A}$ and $A=1,\dots,N_{f}$. The global symmetries include chiral fermion parity $\mathsf{C}$ (sometimes denoted as $(-1)^{F_{\text{L}}}$), reflection $\mathsf{R}$, time-reversal $\mathsf{T}$, and fermion parity $(-1)^{F}$. See (LABEL:eq:contCPT). They form the group $D_{4}\times\mathbb{Z}_{2}$. Here, $D_{4}$ is the dihedral generated by $a=\mathsf{C}\mathsf{R}$ and $b=\mathsf{R}$ with $a^{4}=1$, $b^{2}=1$, and $bab^{-1}=a^{-1}$. The fermion parity operator is $(-1)^{F}=a^{2}$. And the $\mathbb{Z}_{2}$ factor is generated by the anti-unitary transformation $c=\mathsf{R}\mathsf{T}$, which commutes with $a$ and $b$ and $c^{2}=1$. [^37] In Appendix A, we discussed the geometric transformations of space-reflection $\mathscr{R}$ and time-reversal $\mathscr{T}$, and in footnote 31, we noted that $\mathscr{R}\mathscr{T}=(-1)^{F}\mathscr{T}\mathscr{R}$ and $\mathscr{R}^{2}=(-1)^{F}\mathscr{T}^{2}$. As we explained in Appendix A, the field theory transformations $\mathsf{R}$ and $\mathsf{T}$ are combinations of the geometric actions $\mathscr{R}$ and $\mathscr{T}$ with some action on the fields. Therefore, their relations can differ from the geometric relations. As we will see, this group can be realized projectively.

### v2

@@ 4.2 Identifying $\Pi$ that flows to $\Theta$
To show that the symmetry generated by $\Pi$ and $\mathcal{Q}$ is anomaly-free, consider the following Hamiltonian, which is symmetric and trivially gapped: [^25] Following footnote 23, one could ask how we identified $\Pi$ as the operator that flows to $\Theta$, rather than $\Pi_{\text{e}}=\mathcal{X}_{\text{even}}\mathcal{R}\mathcal{T}=X\Pi$, which flows to $\exp(i\pi\mathsf{Q}_{\text{w}})\Theta$. The key point is that the lattice anomaly between $\mathcal{X}$ and $\mathcal{Q}$ leads to a mixed anomaly between $\Pi_{\text{e}}$ and $\mathcal{Q}$. See the discussion in Subsection 4.5.1. On the other hand, there is no anomaly between $\Pi$ and $\mathcal{Q}$ as we show here.
@@ 4.3 Self-anomaly of $\mathcal{U}$
which also probes the self-anomaly of $\mathcal{U}$. This is similar to the previous fermionic and bosonic cases in Subsections 2.3.1 and 3.3.1, where in the presence of “lattice translation defect,” the operator $\tilde{\Pi}$ becomes anomalous. See more comparisons in Section 6.
@@ 4.4 Anomaly for $\tilde{\Pi}$}
Given that there is a mixed anomaly between $\Pi$ and $\mathcal{U}$, it is natural to ask if $\tilde{\Pi}$} itself is anomalous. We will follow the discussion in Section 3.3.1 to identify a mod 2 anomaly for $\tilde{\Pi}$}. We will restrict ourselves to the even $L$ case, and separate the discussion into two cases: $L=0,2$ mod 4.262626In the discussion in Section 4.1, we limited ourselves to $L=0\bmod 4$. This was necessary for the properties of $\mathcal{Q}$. However, here we do not discuss $\mathcal{Q}$, so we can study the model for any even $L$. It will become clear that these two cases are analogous to the even and odd $L$ cases in Section 3.3.1.
@@
[…]
@@ 4.5.1 The lattice anomaly between $\mathcal{X}$ and $e^{i\pi\mathcal{Q}}$
Here, we use $\Pi$ to probe the emergent anomaly between $e^{i\frac{\pi}{2}\mathsf{Q}_{\mathrm{w}}}$ and $e^{i\pi\mathsf{Q}_{\mathrm{m}}}$. (See Section 1.4.)
@@ 5 Non-invertible RT
    $\mathcal{D}X_{j}=Z_{j}Z_{j+1}\mathcal{D}\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \mathcal{D}Z_{j}Z_{j+1}=X_{j+1}\mathcal{D}\,.$ (5.3)
@@
    $\mathcal{T}_{\text{Ising}}X_{j}\mathcal{T}_{\text{Ising}}^{-1}=X_{j}\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \mathcal{T}_{\text{Ising}}Y_{j}\mathcal{T}_{\text{Ising}}^{-1}=-Y_{j}\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \mathcal{T}_{\text{Ising}}Z_{j}\mathcal{T}_{\text{Ising}}^{-1}=Z_{j}\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \mathcal{T}_{\text{Ising}}i\mathcal{T}_{\text{Ising}}^{-1}=-i\,.$ (5.6)
@@
${\tilde{\Pi}}_{\text{Ising}}X_{j}=Z_{-j}Z_{-j+1}{\tilde{\Pi}}_{\text{Ising}}\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ {\tilde{\Pi}}_{\text{Ising}}Z_{j}Z_{j+1}=X_{-j}{\tilde{\Pi}}_{\text{Ising}}\,.$ (5.8)
@@
$({\tilde{\Pi}}_{\text{Ising}})^{2}=1+\eta\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ {\tilde{\Pi}}_{\text{Ising}}\eta=\eta{\tilde{\Pi}}_{\text{Ising}}={\tilde{\Pi}}_{\text{Ising}}\,,\quad\eta^{2}=1\,.$ (5.9)
@@
$i\chi_{2j-1}\chi_{2j}=X_{j}\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ i\chi_{2j}\chi_{2j+1}=Z_{j}Z_{j+1}\,.$ (5.11)
@@ Acknowledgments
We are grateful to Maissam Barkeshli, Boris Bulatovic, Meng Cheng, Lukasz Fidkowski, Tarun Grover, Ryohei Kobayashi, Zohar Komargodski, Ho Tat Lam, Max Metlitski, Greg Moore, Salvatore Pace, Abhinav Prem, Sahand Seifnashri, Nikita Sopenko, Xincheng Zhang, and Yifan Zhang for interesting discussions. The work of NS was supported in part by DOE grant DE-SC0009988. SHS was supported in part by NSF grant PHY-2449936. WZ was supported in part by the Princeton University Department of Physics. This work was also supported by the Simons Collaboration on Ultra-Quantum Matter, which is a grant from the Simons Foundation (651444, NS, SHS).
@@ A.1 Non-spin theories
Let us apply this reasoning to a field theory. As mentioned in footnote 30, now we should use $\mathscr{R}\to\mathsf{R}$, $\mathscr{T}\to\mathsf{T}$, and $\varTheta\to\Theta$. In addition to the invariance under $\Theta$, the theory might also have additional symmetries. These can be internal symmetries, which do not act on the spacetime coordinate, as well as spacetime symmetries like space-reflection and time-reversal. We would like to make a few comments about the consequences of that.
@@ A.2 Spin theories
In Euclidean signature, we lift $SO(d+1)$ to $Spin(d+1)$ and again $R_{ij}(2\pi)=(-1)^{F}$ with $i,j=1,\cdots,d+1$ and $i\neq j$. [^31] Even though we do not focus on the individual reflection elements $\mathscr{R}^{i}$, but only on their bilinears $\mathscr{R}^{i}\mathscr{R}^{j}=R_{ij}(\pi)$, we note that $R_{ij}(\pi)=(-1)^{F}R_{ji}(\pi)$. And therefore, for $i\neq j$, $\mathscr{R}^{i}\mathscr{R}^{j}=(-1)^{F}\mathscr{R}^{j}\mathscr{R}^{i}$. (This is regardless of whether $(\mathscr{R}^{i})^{2}$ is $1$ or $(-1)^{F}$, i.e., regardless of whether they extend $Spin(d+1)$ to $Pin^{+}(d+1)$ or to $Pin^{-}(d+1)$.) This means that after the lift $R_{ij}(\pi)^{2}=(-1)^{F}$, including
@@
Rotating this back to Lorentzian signature, $\mathscr{R}^{d+1}$ becomes the anti-unitary transformation $\mathscr{T}$. Therefore, $\mathscr{R}\mathscr{R}^{d+1}$ becomes the anti-unitary transformation $\varTheta=\mathscr{R}\mathscr{T}$ and its square has another factor of $(-1)^{F}$ to give [^32] As in footnote 31, $\mathscr{R}\mathscr{T}=(-1)^{F}\mathscr{T}\mathscr{R}$, and using (A.16), $\mathscr{R}^{2}=(-1)^{F}\mathscr{T}^{2}$. Also, $\Theta\mathscr{T}=(-1)^{F}\mathscr{T}\Theta$, $\Theta\mathscr{R}=(-1)^{F}\mathscr{R}\Theta$, and $\Theta\mathscr{R}^{i}=\mathscr{R}^{i}\Theta$ for $i\neq 1$. (Again, this is regardless of whether $\mathscr{R}^{2}$ is $1$ or $(-1)^{F}$.)
@@
This crucial relation will be discussed in detail below for the special case of $d=1$. The extension to $d>1$ is straightforward.
@@
We claim that a Hermitian operator $\cal O$ with spin $s$ has [^34] The anti-unitary operator $\varTheta$ acts on a local operator $\mathcal{O}(t,x)$ as $\varTheta\mathcal{O}(t,x)\varTheta^{-1}=\xi\mathcal{O}(-t,-x)^{\dagger}$ with a phase $\xi$. (We have suppressed possible Lorentz indices of the local operator.) Note that the phase $\xi$ is invariant under operator redefinitions because of the adjoint operation on the right-hand side. To define the eigenvalues of $\varTheta$, we need to further assume that the operator $\mathcal{O}$ is Hermitian. This fact is at the root of the $\varTheta$ assignment for spinors in (LABEL:Rof_O). In this case we denote the eigenvalue as $\xi=\varTheta(\mathcal{O})$.
@@
In order to understand the significance of the assignment of $\varTheta({\cal O})$, let us discuss the multiplication of these representations. Consider two Hermitian operators, ${\cal O}_{i}$ with spins $s_{i}$ and $\varTheta$ and $(-1)^{F}$ as in (LABEL:Rof_O). If $s_{1}$ is an integer, ${\cal O}_{1}{\cal O}_{2}$ is Hermitian. Its total spin is $s_{1}+s_{2}$ and $\varTheta({\cal O}_{1}{\cal O}_{2})=(-1)^{s_{1}}(-1)^{s_{2}(2s_{2}-1)}=(-1)^{(s_{1}+s_{2})(2s_{1}+2s_{2}-1)}$. The same conclusion applies if $s_{2}$ is an integer. However, if both $s_{1}$ and $s_{2}$ are half-integers, these two operators are fermions. Then, to determine the $\varTheta$-eigenvalue of their product, we should consider the Hermitian operator $i{\cal O}_{1}{\cal O}_{2}$. (See footnote 34.) Its spin is $s_{1}+s_{2}$ and $\varTheta(i{\cal O}_{1}{\cal O}_{2})=-(-1)^{s_{1}(2s_{1}-1)}(-1)^{s_{2}(2s_{2}-1)}=(-1)^{s_{1}+s_{2}}$. (This reasoning provides a derivation of the $\varTheta$-eigenvalue assignments in (LABEL:Rof_O) by multiplying representations with $s=\pm{1\over 2}$.)
@@
The individual transformations $\mathsf{C}$, $\mathsf{R}$, and $\mathsf{T}$ might not be symmetries, and even if they are, they are subject to redefinitions using internal symmetries. However, regardless of such redefinitions, when they exist,
$\mathsf{T}\Theta=(-1)^{F}\Theta\mathsf{T}\quad,\quad\mathsf{R}\Theta=(-1)^{F}\Theta\mathsf{R}\quad,\quad\mathsf{R}^{i}\Theta=\Theta\mathsf{R}^{i}\qquad{\rm for}\ i\neq 1\,.$ (A.24)
where $\mathsf{R}^{i}$ acts as $\mathsf{R}$ except that it reflects $x^{i}$. Geometrically, this follows from the expressions in footnote 32. (Intuitively, for $i\neq 1$, the symmetry $\mathsf{R}^{i}$ acts as an internal symmetry as far as $x^{1}$ and $t$ are concerned, and therefore, as in (1.3), it commutes with $\Theta$.) For example, these relations can be verified explicitly in the example of the 1+1d Majorana fermion by using (LABEL:eq:contCPT). Note that the geometric relation in footnote 32 $\mathscr{R}^{2}=(-1)^{F}\mathscr{T}^{2}$, might not lead to a similar relation for $\mathsf{R}$ and $\mathsf{T}$ because their redefinitions using internal transformations might change it. Indeed, in (LABEL:eq:contCPT), we have $\mathsf{R}^{2}=\mathsf{T}^{2}=1$.
Finally, repeating the discussion about odd spatial dimensions in (A.13), we have
@@
And if we have $\mathsf{R}$ and $\mathsf{T}$ symmetries, (A.24) is replaced with
$\mathsf{T}(\mathsf{C}\mathsf{P}\mathsf{T})=(-1)^{F}(\mathsf{C}\mathsf{P}\mathsf{T})\mathsf{T}\quad,\quad\mathsf{R}^{i}(\mathsf{C}\mathsf{P}\mathsf{T})=(-1)^{F}(\mathsf{C}\mathsf{P}\mathsf{T})\mathsf{R}^{i}\quad,\quad\forall i\,.$ (A.26)
@@ Appendix B Review of time-reversal anomalies in quantum mechanics of fermions
$[(-1)^{F}]^{2}=1\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \mathcal{T}^{2}=(-1)^{\frac{\nu(\nu-2)}{8}}\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \mathcal{T}(-1)^{F}=(-1)^{\frac{\nu}{2}}(-1)^{F}\mathcal{T}\,,\quad\nu=0\bmod 2\,.$ (B.10)
@@
$\mathcal{T}^{\prime 2}=(-1)^{\frac{\nu(\nu+2)}{8}}\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \mathcal{T}^{\prime}(-1)^{F}=(-1)^{\frac{\nu}{2}}(-1)^{F}\mathcal{T}^{\prime}\,,\quad\nu=0\bmod 2\,.$ (B.11)
@@
We consider $N_{f}$ non-chiral, massless, free Majorana fermion fields in the continuum (2.16), with their left- and right-moving components denoted by $\chi_{\text{L,R}}^{A}$ and $A=1,\dots,N_{f}$. The global symmetries include chiral fermion parity $\mathsf{C}$ (sometimes denoted as $(-1)^{F_{\text{L}}}$), reflection $\mathsf{R}$, time-reversal $\mathsf{T}$, and fermion parity $(-1)^{F}$. See (LABEL:eq:contCPT). They form the group $D_{4}\times\mathbb{Z}_{2}$. Here, $D_{4}$ is the dihedral generated by $a=\mathsf{C}\mathsf{R}$ and $b=\mathsf{R}$ with $a^{4}=1$, $b^{2}=1$, and $bab^{-1}=a^{-1}$. The fermion parity operator is $(-1)^{F}=a^{2}$. And the $\mathbb{Z}_{2}$ factor is generated by the anti-unitary transformation $c=\mathsf{R}\mathsf{T}$, which commutes with $a$ and $b$ and $c^{2}=1$. [^38] In Appendix A, we discussed the geometric transformations of space-reflection $\mathscr{R}$ and time-reversal $\mathscr{T}$, and in footnote 32, we noted that $\mathscr{R}\mathscr{T}=(-1)^{F}\mathscr{T}\mathscr{R}$ and $\mathscr{R}^{2}=(-1)^{F}\mathscr{T}^{2}$. As we explained in Appendix A, the field theory transformations $\mathsf{R}$ and $\mathsf{T}$ are combinations of the geometric actions $\mathscr{R}$ and $\mathscr{T}$ with some action on the fields. Therefore, their relations can differ from the geometric relations. As we will see, this group can be realized projectively.